# Single-Shot Decoding and Fault-tolerant Gates with Trivariate Tricycle Codes

### v1

@@ III Single-shot decoding of TT codes under phenomenological noise
In Fig. 2(a), we present data for the $X$ memory experiments, examining the codes presented in Table 1. We plot logical error probability (per round) for each code, without normalising for the number of logical qubits. We choose to run the memory experiment simulations for $N=2d$ rounds, where $d$ is the code distance. Decoding is performed using a BP+OSD decoder (without any windowing) [47, 48] and we perform quadratic fits to the data. We additionally plot the results for the $[[81,3,3]]$ 3DTC for comparison, showing that all TT codes tested out-perform this code for the range of physical error rates examined.
@@
We also examine the $Z$ memory performance under different windowing strategies. In Fig. 3 we present results for a $(w,1)$ windowing strategy, where we vary $w$ between 1 and 6. We present results for a range of codes from Table 1 and Table 5. We fix the total number of syndrome extraction rounds to be $N=14$ and the physical error rate to be $p=3\times 10^{-2}$. We find that the logical error rates for $Z$ memory experiments for the TT codes rapidly plateau as we increase the window size, meaning that the performance of the decoder does not continue to increase as more syndrome information is taken into account during each round of decoding, indicating single-shot properties.
@@ IV Circuit Level Noise Simulations
We conducted $X$ and $Z$ memory simulations, and verified the existence of $(w,c)$ single-shot decodability under a circuit level noise model for both the TT and 3D Toric codes. Our simulations were carried out using the Python packages LDPC [48], Stim [52] and stimbposd [53]. We utilise depth-13, interleaved syndrome measurement circuits detailed in Appendix B. We omit resets on the $X$-measure qubits to minimise the circuit depth, and note that this likely has a minimal effect on memory performance [54, 55]. We define three sets of $Z$-measure qubits (for measuring $Z_{a}$, $Z_{b}$ and $Z_{c}$ checks), and for two of these, we alternate using them for $Z_{b}$ and $Z_{c}$ checks in adjacent rounds, again to minimise circuit-depth. Overall we use $4n/3$ measure qubits (for $n$ data qubits in the code). In Appendix B, we prove that the circuits result in the required transformations of the stabilizer tableaux. We stress that we have not optimised the ordering of these circuits for circuit-level distance, and further improvements could potentially be made.
@@
where $P_{L}$ is the total number of logical errors over all shots. To extrapolate the code performance to lower error rates, we fit trend lines using the ansatz, $p_{L}=p^{d_{circ}/2}e^{c_{0}+c_{1}p+c_{2}p^{2}}$ [12], and numerically compute the coefficients using SciPy [56]. The parameter $d_{circ}$ is the upper-bound on the circuit level distance, obtained by finding the minimum weight error that triggers no detectors in the detector matrix [57]. The detector matrix is obtained using stimbposd and the minimum weight error was found using the technique in Section 6 of Ref. [12] and the BP+OSD implementation in Ref. [48]. For large detector matrices, for example in the $[[432,12,12]]$ code, BP+OSD struggles to find low weight errors when estimating circuit-level distance. In that case, in fits we use an effective circuit distance $d^{\prime}_{circ}\leq d_{circ}$, the maximum value that produces a trend line that fits the data points. Due to the asymmetry in the $X$ and $Z$ distances, the limiting distance is the $Z$ distance. As such, in Table 2 we present crossing points for the $X$ memory experiments, which serve as an estimate for a threshold for the TT codes. In the following, we do not normalise for the number of logical qubits encoded, hence the error rates include correlated errors within a code patch.
@@ IV.1 Simulation Results
We perform $Z$ memory, single-shot experiments under a $(w,1)$ strategy for various values of $w$ and show that a logical error plateau is recovered even under a circuit level noise model. This suggests that TT codes retain single-shot capabilities in realistic settings. In Fig.5 we explore sub-threshold noise regimes while varying the window size. We observe that the code performance rapidly plateaus as we increase the window size, albeit at a slower rate for larger $p$. That is, at a physical error rate of $p=6\times 10^{-3}$ the LER plateau occurs at around $w=5$, whereas at a physical error rate of $p=3\times 10^{-3}$ the LER plateau begins at around $w=3$. It is also interesting to note that the $[[72,6,6]]$ codes has a higher logical error rate than the $[[81,6,6]]$ despite having the same $X$-distance and fewer qubits. We suspect this discrepancy is caused by structural differences in the Tanner graphs of the two codes. Of relevance could be the fact that the $[[81,6,6]]$ code has a toric layout, while the $[[72,6,6]]$ code does not. A thorough analysis is deferred to future work.
@@ V.3 Constant-depth $CCZ$ circuits
We finally consider the possibility of fault-tolerant non-Clifford gates in TT codes. On the 3D toric code on a cubic lattice, one can implement a set of logical $\overline{CCZ}$ gates via a depth-2 circuit of physical $CCZ$ gates on the data qubits using a construction [62, 63, 64] derived from the cohomological cup product [35]. Similarly, any TT code whose checks are based on weight-2 polynomials, $A=A_{1}+A_{2}$, $B=B_{1}+B_{2}$, and $C=C_{1}+C_{2}$, admits a logical $\overline{CCZ}$ implementable in the same way – we explain this in the next section. We will call such a code a (2,2,2) TT code, with the numbers referring to the weights of the polynomials $A$, BB and $C$, respectively.
@@ V.3.1 TT Codes with Cup Products
Table 4: Examples of (4,2,2) codes with CCZCCZ gates that have non-trivial logical action. Third column: Rather than simply reporting the $Z$-distance (which is 2 in all cases), we report the minimum weights found for all independent $Z$-logical operators. Fifth column: The AA polynomial is chosen to have weight-4, and to be of the form $(1+g)(1+a)$, where $g^{2}=1$ and $a\neq 1$ is arbitrary. Sixth column: Number of logical $\overline{CCZ}$ gates implemented by the constant-depth circuit of physical $CCZ$ gates in a basis defined by the $Z$-logical operators with weights given in the third column.
@@ Appendix B Syndrome measurement circuits
Note that these circuits have the complication that the $Z_{b}$ and $Z_{c}$ blocks of measure qubits switch labels every round – this is indicated after the resets in Fig. 7. In order to not introduce an extra idling layer, we omit explicit resets on the $X$-measure qubits, instead performing a Pauli frame update dependent on the outcomes of measurements of the $X$-measure qubits (alternatively, one could assume having an additional $n/3$ $X$-measure qubits available). The lack of resets can be accommodated by updating the definitions of detectors (see, e.g., Ref. [54]). This should not drastically affect quantum memory performance, though it decreases the time-like distance of certain logical operations, like (generalised) lattice surgery (see, e.g., Ref. [54, 55]). For simplicity, in our simulations, we perform a noiseless reset rather than update the definitions of detectors. As drawn, the circuits have depth 13. The first reset in a layer is on the $Z_{a}$-measure qubits and the final measurement is on the $X$-measure qubits. If we perform resets on the $X$-measure qubits, the circuits become depth-14, however the 2-qubit gate depth remains the same.
@@ C.2 Proof of Lemma 5
It suffices to show that the polynomial $P$ satisfies the cup product conditions, since $A$, BB and $C$ must do so individually, and we know that weight-2 polynomials satisfy the conditions.
@@ C.4 Details of numerics used to find logical action of $CCZ$ circuits
Input: Matrices $H_{X}$, $H_{Z}$; List of $CCZ$ gates in circuit 99, $L$
@@
for each $\texttt{row}_{1}$ in rs($H_{X}$) do
 for each $\texttt{row}_{2}$ in rs($H_{X}$) do
    $Z$-stabilizer candidate $\leftarrow$ row vector of zeros with same number of columns as $H_{Z}$;
    for each $\texttt{qubit}_{1}$ in $\texttt{row}_{1}$ and $\texttt{qubit}_{2}$ in $\texttt{row}_{2}$ do
       if $\texttt{qubit}_{1}$ and $\texttt{qubit}_{2}$ are in support of the same CCZCCZ gate $G$ then
            Third qubit $\leftarrow$ row vector with 1 in column corresponding to the third qubit acted upon by $G$, and zeros in all other columns;
          $Z$-stabilizer candidate $\leftarrow$ $Z$-stabilizer candidate ++ Third qubit;

### v2

@@ III Single-shot decoding of TT codes under phenomenological noise
In Fig. 2(a), we present data for the $X$ memory experiments, examining the codes presented in Table 1. We plot logical error probability (per round) for each code, without normalising for the number of logical qubits. We choose to run the memory experiment simulations for $r=2d$ rounds, where $d$ is the code distance. Decoding is performed using a BP+OSD decoder (without any windowing) [51, 52] and we perform quadratic fits to the data. We additionally plot the results for the $[[81,3,3]]$ 3DTC for comparison, showing that all TT codes tested out-perform this code for the range of physical error rates examined.
@@
We also examine the $Z$ memory performance under different windowing strategies. In Fig. 3 we present results for a $(w,1)$ windowing strategy, where we vary $w$ between 1 and 6. We present results for a range of codes from Table 1 and Table 5. We fix the total number of syndrome extraction rounds to be $r=14$ and the physical error rate to be $p=3\times 10^{-2}$. We find that the logical error rates for $Z$ memory experiments for the TT codes rapidly plateau as we increase the window size, meaning that the performance of the decoder does not continue to increase as more syndrome information is taken into account during each round of decoding, indicating single-shot properties.
@@ IV Circuit Level Noise Simulations
We conduct $X$ and $Z$ memory simulations, and verify the existence of $(w,c)$ single-shot decodability under a circuit level noise model for both the TT and 3D Toric codes. We utilise depth-13, interleaved syndrome measurement circuits detailed in Appendix B. We define three sets of $n/3$ $Z$-measure qubits (for measuring $Z_{a}$, $Z_{b}$ and $Z_{c}$ checks), and for two of these, we alternate using them for $Z_{b}$ and $Z_{c}$ checks in adjacent rounds, again to minimise circuit-depth, lastly we use $n/3$ $X$-measure qubits. However, we note that omitting resets on the $X$-measure qubits to minimise the circuit depth, likely has a minimal effect on memory performance [gehér2025resetresetquestion, 25]. Overall we use $4n/3$ measure qubits (for $n$ data qubits in the code). In Appendix B, we prove that the circuits result in the required transformations of the stabilizer tableaux. We stress that we have not optimised the ordering of these circuits for circuit-level distance, and further improvements could potentially be made.
@@
where $P_{L}$ is the total number of logical errors over all shots. To extrapolate the code performance to lower error rates, we fit trend lines using the ansatz, $p_{L}=p^{d_{circ}/2}e^{c_{0}+c_{1}p+c_{2}p^{2}}$ [5], and numerically compute the coefficients using SciPy [59]. The parameter $d_{circ}$ is the upper-bound on the circuit level distance, obtained by finding the minimum weight error that triggers no detectors in the detector matrix [17]. The detector matrix is obtained using stimbposd [29] and the minimum weight error was found using the technique in Section 6 of Ref. [5] and BP+OSD [52]. For large detector matrices, for example in the $[[432,12,12]]$ code, BP+OSD struggles to find low weight errors when estimating circuit-level distance. In that case, in fits we use an effective circuit distance $d^{\prime}_{circ}\leq d_{circ}$, the maximum value that produces a trend line that fits the data points. Due to the asymmetry in the $X$ and $Z$ distances, the limiting distance is the $Z$ distance. As such, in Table 2 we present crossing points for the $X$ memory experiments, which serve as an estimate for a threshold for the TT codes. In the following, we do not normalise for the number of logical qubits encoded, hence the error rates include correlated errors within a code patch.
@@ IV.1 Simulation Results
We perform $Z$ memory, single-shot experiments under a $(w,1)$ strategy for various values of $w$ and show that a logical error plateau is recovered even under a circuit level noise model. This suggests that TT codes retain single-shot capabilities in realistic settings. In Fig.5 we explore sub-threshold noise regimes while varying the window size. We observe that the code performance rapidly plateaus as we increase the window size, albeit at a slower rate for larger $p$. That is, at a physical error rate of $p=6\times 10^{-3}$ the LER plateau occurs at around $w=5$, whereas at a physical error rate of $p=3\times 10^{-3}$ the LER plateau begins at around $w=3$. It is also interesting to note that the $[[72,6,6]]$ code has a higher logical error rate than the $[[81,6,6]]$ despite having the same $X$-distance and fewer qubits. We suspect this discrepancy is caused by structural differences in the Tanner graphs of the two codes. Of relevance could be the fact that the $[[81,6,6]]$ code has a toric layout, while the $[[72,6,6]]$ code does not. A thorough analysis is deferred to future work.
@@ V.3 Constant-depth $CCZ$ circuits
We finally investigate the possibility of fault-tolerant non-Clifford gates in TT codes. Fault-tolerant logical CCZ gates have been implemented on certain QLDPC codes using a construction derived from the cohomological cup product  [69, 14, 63, 9]. This technique has recently been applied to hypergraph product codes and homological products of LDPC codes, showing favourable asymptotic scaling [70, 71].
We show below that any TT code whose checks are based on weight-2 polynomials, $A=A_{1}+A_{2}$, $B=B_{1}+B_{2}$, and $C=C_{1}+C_{2}$, admits a logical $\overline{CCZ}$. Crucially, this gate is implementable via a depth-2 physical CCZ circuit, similar to what is used for the 3D Toric code on a cubic lattice. We call such a code a (2,2,2) TT code, with the numbers referring to the weights of the polynomials A, B and C, respectively; we detail this construction in the next section.
@@ V.3.1 TT Codes with Cup Products
Table 4: Examples of (4,2,2) codes with CCZCCZ gates that have non-trivial logical action. Third column: Rather than simply reporting the $Z$-distance (which is 2 in all cases), we report the minimum weights found for all independent $Z$-logical operators. Fifth column: The AA polynomial is chosen to have weight-4, and to be of the form $(1+g)(1+a)$, where $g^{2}=1$ and $a\neq 1$ is arbitrary. Sixth column: Number of logical $\overline{CCZ}$ gates implemented by the constant-depth circuit of physical CCZCCZ gates in a basis defined by the $Z$-logical operators with weights given in the third column.
@@ Appendix B Syndrome measurement circuits
Note that these circuits have the complication that the $Z_{b}$ and $Z_{c}$ blocks of measure qubits switch labels every round – this is indicated after the resets in Fig. 7. In order to not introduce an extra idling layer, we omit explicit resets on the $X$-measure qubits, instead performing a Pauli frame update dependent on the outcomes of measurements of the $X$-measure qubits (alternatively, one could assume having an additional $n/3$ $X$-measure qubits available). The lack of resets can be accommodated by updating the definitions of detectors (see, e.g., Ref. [gehér2025resetresetquestion]). This should not drastically affect quantum memory performance, though it decreases the time-like distance of certain logical operations, like (generalised) lattice surgery (see, e.g., Ref. [gehér2025resetresetquestion, 25]). For simplicity, in our simulations, we perform a noiseless reset rather than update the definitions of detectors. As drawn, the circuits have depth 13. The first reset in a layer is on the $Z_{a}$-measure qubits and the final measurement is on the $X$-measure qubits. If we perform resets on the $X$-measure qubits, the circuits become depth-14, however the 2-qubit gate depth remains the same.
@@ C.2 Proof of Lemma 5
It suffices to show that the polynomials of weight 2 or of the form of $P$ satisfy the cup product conditions, since $A$, BB and $C$ must satisfy these conditions individually. To see that a weight-2 polynomial, $P_{1}+P_{2}$, satisfies the conditions, simply let $P_{\text{in}}\coloneqq P_{1}$ and $P_{\text{out}}\coloneqq P_{2}$. All conditions are thereby easily verified.
@@ C.4 Details of numerics used to find logical action of $CCZ$ circuits
The intuition for the algorithm is that it checks whether the $CCZ$ circuit conjugates $X$-stabilizers in any two code blocks to a $Z$-stabilizer in the third block. We note in the algorithm that rows(A)(A) corresponds to the set of rows of matrix $A$, while rs(A)(A) corresponds to the row space of $A$ (i.e., the span of all rows of $A$ over the field $\mathbb{F}_{2}$). We also note that it suffices to pick one stabilizer (labelled $\texttt{row}_{1}$ in Algorithm 1) of the first code block arbitrarily before looping over only the second-code-block $X$-stabilizer ($\texttt{row}_{2}$). This is because all other first-block $X$-stabilizers are related to $\texttt{row}_{1}$ by a monomial. This monomial can be absorbed into the product over $\alpha\in\mathcal{M}$ in the circuit (Equation 99) and hence if the circuit conjugates $\texttt{row}_{1}$ and $\texttt{row}_{2}$ to a $Z$-stabilizer (for all choices of $\texttt{row}_{2}$) it does so for all choices of $\texttt{row}_{1}$. This is simply a result of the translation invariance of the code.
@@
Input: Matrices $H_{X}$, $H_{Z}$; List of $CCZ$ gates in circuit 99, $L$ (each is a length-3 tuple of qubit indices corresponding to the support of that gate in the three code blocks)
@@
for $i$ in $(1,2,3)$ and j in $(1,2,3)$ such that $i\neq j$ do
 $\texttt{row}_{1}\leftarrow$ First row of $H_{X}$;
   Ind($\texttt{row}_{1}$) $\leftarrow$ List of indices corresponding to non-zero elements of $\texttt{row}_{1}$;
 $n\leftarrow$ number of columns in $H_{X}$;
 for each $\texttt{row}_{2}$ in rows($H_{X}$) do
      Ind($\texttt{row}_{2}$) $\leftarrow$ List of indices corresponding to non-zero elements of $\texttt{row}_{2}$;
    $Z$-stabilizer candidate $\leftarrow$ row vector of zeros with length $n$;
    for each $\texttt{qubit}_{1}$ in Ind($\texttt{row}_{1}$) and $\texttt{qubit}_{2}$ in Ind($\texttt{row}_{2}$) do
       if $\texttt{qubit}_{1}$ and $\texttt{qubit}_{2}$ are in support of the same CCZCCZ gate $G\in L$, with $\texttt{qubit}_{1}$ ($\texttt{qubit}_{2}$) in the $i^{\text{th}}$ ($j^{\text{th}}$) argument of $G$ then
          qubitvec $\leftarrow$ row vector of length $n$ with 1 in the column corresponding to the qubit in the $k^{\text{th}}$ argument of $G$ (where $k\neq i,k\neq j$), and zeros in all other columns;
          $Z$-stabilizer candidate $\leftarrow$ $Z$-stabilizer candidate ++ qubitvec;